OpenStudy (anonymous):

a factory can produce two product, x and y, with a profit approximated by p=14x+22y-900. the production of y can exceed x by no more than 100 units. moreover, productions levels are limited by the formula x+2y is less than or equal two 1,400. what production levels yield maximum profit?

OpenStudy (mathstudent55):

These are the constraints: p=14x+22y-900 y - x <= 100 x + 2y <= 1400

OpenStudy (anonymous):

yeah

OpenStudy (mathstudent55):

The production of x and y must each be 0 or more. The constraints we have are: \(x \ge 0\) \(y \ge 0\) \(y - x \le 100\) \(x + 2y \le 1400\) Now we can look at all the corresponding equations and graph them to find a region. \(x = 0\) \(y = 0\) \(y - x = 100 \rightarrow y = x + 100\) \(x + 2y = 1400 \rightarrow y = -\dfrac{1}{2}x + 700\)

OpenStudy (mathstudent55):